OpenStudy (anonymous):

Which factors determine the strength of this electromagnet? A. the amount of electric current and the number of wire coils around the nail B. the number of coils of wire around the nail and the length of the wire in the circuit C. the size of the nail and the size of the switch. D. the voltage of the battery and the size of the switch

OpenStudy (anonymous):

is it B?

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

 According to Ampere's Law $$ \int\vec B\cdot d\vec{l}=B\times L=\mu_0NI\\ B=\cfrac{\mu_0NI}{L}~Tesla $$ Where N=number of windings, I is current, L is the length of the coil. Does this help?